OpenStudy (anonymous):

find the i and j component of v in terms of the magnitude of v ad theta simplifying your answer (theta is in the second quadrant )

OpenStudy (anonymous):

can somebody hlp me please

OpenStudy (anonymous):



OpenStudy (anonymous):

i = v*cos(180-theta) j = v*sin(180-theta)

OpenStudy (anonymous):

Thanks a lot that's what I got but couldn't simplify it much more

OpenStudy (anonymous):

great! It's already simplified. Note that in order to obtain vector components, let's say along the x-axis or i, you have to start at 0 degrees. but since the given angle theta is at the second quadrant, I subtracted 180 to theta to get the angle starting from zero.  with that, we can say that: i = v*cos(180-theta) = v*cos(alpha) j = v*sin(180-theta) = v*sin(alpha)